How many orbitals and electrons are associated with n=4?

A

32, 64

B

16, 32

C

4, 16

D

8, 16

 

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
Fornthorbital possible values of azimuthal quantum number (subshell), l are from 0 to (n-1). Total of 'n' values.
Inn=4,l=0,1,2,3thus there are 4 subshells i.e.s,p,d,frespectively.
Magnetic quantum numberml​have values from -l to +l and total of2l+1values.
Forn=4, possible values oflandml​are:
ml​=0for l=0; totalml​values =1
ml​=−1,0,1for l=1; totalml​values =3
ml​=−2,−1,0,1,2for l=2; totalml​values =5
ml​=−3,−2,−1,0,1,2,3for l=3; totalml​values =7
Total number of orbitals = total values ofml​
for n=4,
∴1+3+5+7=16orbitals
Each orbital can occupy maximum of two electron
Number of electrons=2×16=32
Therefore in n=4, number of subshells=4, orbitals=16 and number of electrons =32.